OpenStudy (anonymous):

Find all polar coordinates of point P where P =(1,-π/6)

OpenStudy (anonymous):

@Mertsj Can you help?

OpenStudy (anonymous):

@ParthKohli please help I can't figure it out

OpenStudy (mertsj):



OpenStudy (mertsj):

\[(1,\frac{11\pi}{6}) + 2\pi n\] \[(1,-\frac{\pi}{6})+2\pi n\]

OpenStudy (anonymous):

Sorry to be a pain in your butt but I don't 100% understand what answer it is :( A. (1, -π/6 + (2n + 1)π) or (-1, -π/6+ 2nπ) B. (1, -π/6 + 2nπ) or (-1, -π/6 + 2nπ) C. (1, -π/6 + 2nπ) or (1, π/6 + (2n + 1)π) D. (1, -π/6 + 2nπ) or (-1, -π/6 + (2n + 1)π) @Mertsj

OpenStudy (mertsj):

Oh yes. I forgot the -1 possibility for the radius

OpenStudy (mertsj):

Are you sure you posted those answers exactly right?

OpenStudy (mertsj):